OpenStudy (anonymous):

Solve. Show your work 1/2 x + 1/4 = 5/16

OpenStudy (anonymous):

0.5x + 0.25 = (5/16) -0.25 -0.25 0.5x = (1/16) divide both sides by 0.5 x = (1/8) (1/2)(1/8)+(1/4)=(5/16)? (1/16)+(1/4)=(5/16)? (5/16) = (5/16)

OpenStudy (anonymous):

thank you